# Time-MMD: A New Multi-Domain Multimodal Dataset for Time Series Analysis

### v1

@@ Abstract
Time series data are ubiquitous across a wide range of real-world domains. While real-world time series analysis (TSA) requires human experts to integrate numerical series data with multimodal domain-specific knowledge, most existing TSA models rely solely on numerical data, overlooking the significance of information beyond numerical series. This oversight is due to the untapped potential of textual series data and the absence of a comprehensive, high-quality multimodal dataset. To overcome this obstacle, we introduce Time-MMD, the first multi-domain, multimodal time series dataset covering 9 primary data domains. Time-MMD ensures fine-grained modality alignment, eliminates data contamination, and provides high usability. Additionally, we develop MM-TSFlib, the first multimodal time-series forecasting (TSF) library, seamlessly pipelining multimodal TSF evaluations based on Time-MMD for in-depth analyses. Extensive experiments conducted on Time-MMD through MM-TSFlib demonstrate significant performance enhancements by extending unimodal TSF to multimodality, evidenced by over 15% mean squared error reduction in general, and up to 40% in domains with rich textual data. More importantly, our datasets and library revolutionize broader applications, impacts, research topics to advance TSA. The dataset and library are available at https://github.com/AdityaLab/Time-MMD and https://github.com/AdityaLab/MM-TSFlib.
@@ 1 Introduction
Time series (TS) data are ubiquitous across a wide range of domains, including economics, urban computing, and epidemiology (33; 36; 19). Analytical tasks on such datasets hence find broad applications in various real-world scenarios such as energy forecasting, traffic planning, and epidemic policy formulation. Human experts typically complete such Time-Series Analysis (TSA) tasks by integrating multiple modalities of time-series data. For instance, epidemiologists combine numerical data on influenza infections with textual domain knowledge, policies, and reports to predict future epidemiological trends. However, most existing TSA models (38; 26; 54; 28; 56; 20; 55; 30) are unimodal, solely using numerical series.
Recently, with the development of Large Language Models (LLMs), the field of TSA is also undergoing an exciting transformative moment with the integration of natural language (50; 21). Existing LLM-based TSA methods incorporate endogenous text derived from numerical series, such as linguistic descriptions of statistical information, which has demonstrated promising benefits (15; 18; 4; 24). However, the potential of exogenous or auxiliary textual signals—such as information on concurrent events and policies that provide additional context to time series—remains untapped. This observation prompts a crucial question for multimodal TSA: Can multimodal TSA models utilize these exogenous textual signals effectively, thereby enhancing current TSA tasks and enabling new applications?
The primary obstacle in addressing this question lies in the absence of a comprehensive, high-quality multimodal TS dataset, as evidenced by three significant gaps: (1) Narrow data domains. Data characteristics and patterns vary between different domains, such as the periodicity of numerical data and the sparsity of textual data. However, current multimodal TS datasets (12; 48; 9; 31; 4) focus solely on stock prediction tasks in the financial domain, which are unable to represent the diverse data domains. (2) Coarse-grained modality alignment. Existing multimodal TS datasets only ensure that the text and numerical data come from the same domain, such as general stock news and the prices of one specific stock. Clearly, an abundance of irrelevant text diminishes the effectiveness of multimodal TSA. (3) Inherent data contamination. Existing multimodal TS datasets overlook two main reasons of data contamination: (1) Textual data often contains predictions. For example, influenza outlook is a regular section in influenza reports. (2) Outdated test set, particularly the textual data, may have been exposed to LLMs, which are pretrained on vast corpuses. For example, the knowledge cutoff for Llama3-70B is December 2023, which is later than the cutoff dates for most existing multimodal TS datasets. These reasons lead to biased evaluations of general or LLM-based TSA models.
@@
Pioneering Multi-Domain Multimodal Time-Series Dataset: Time-MMD. We introduce Time-MMD, the first multi-domain multimodal time-series dataset that addresses the aforementioned gaps: (1) encompasses 9 primary data domains. (2) ensures fine-grained modality alignment through meticulously selected data sources and rigorous filtering steps. (3) disentangles facts and predictions from text; ensures all cutoff dates are up to May 2024. To the best of our knowledge, Time-MMD stands as the inaugural high-quality and comprehensive multimodal time-series dataset. We envision Time-MMD offering exciting opportunities to significantly advance time series analysis through multimodal extensions.
Pilot Multimodal Time-Series Forecasting Library: MM-TSFlib. We develop the first multimodal time-series forecasting (TSF) library, MM-TSFlib, piloting multimodal TSA research based on Time-MMD. Our library MM-TSFlib features an end-to-end pipeline with a seamless interface that allows the integration of any open-source language models with arbitrary TSF models, thereby enabling multimodal TSF tasks. MM-TSFlib facilitates easy exploration of Time-MMD and supports future advancements in multimodal TSA.
Extensive Evaluations with Significant Improvement. We conducted experiments of multimodal TSF on Time-MMD using MM-TSFlib. The multimodal versions outperformed corresponding unimodal versions in all TSF backbones, reducing the mean squared error by an average of over 15% and up to 40% in some domains with rich textual data. This significant and consistent improvement demonstrates the high quality of Time-MMD, the effectiveness of MM-TSFlib, and the superiority of multimodal extensions for TSF.
@@ 2.1 Numerical Series Data Construction
Data Source Selection. We select data sources that are (1) reliable, containing verified knowledge; (2) actively released, allowing for updates with new data; and (3) multi-domain, covering various TSA patterns. Based on these principles, we choose nine data sources from different domains. Most sources are from government agencies, with the lowest update frequency being six months.
@@
| Domain | Target | Frequency | Timestamps | Timespan |
| Agriculture | Retail Broiler Composite | Monthly | 1983 - Present | |
| Climate | Drought Level | Monthly | 1983 - Present | |
| Economy | International Trade Balance | Monthly | 1989 - Present | |
| Energy | Gasoline Prices | Weekly | 1996 - Present | |
| Environment | Air Quality Index | Daily | 1982 - 2023 | |
| Health | Influenza Patients Proportion | Weekly | 1997 - Present | |
| Security | Disaster and Emergency Grants | Monthly | 1999 - Present | |
| Social Good | Unemployment Rate | Monthly | 1950 - Present | |
| Traffic | Travel Volulmn | Monthly | 1980 - Present |
@@ 2.2 Textual Series Data Construction
These two data sources complement each other: report data ensures higher relevance but cannot guarantee all-time coverage, while search results cover all times but are highly redundant; search results aggregate multiple data sources, while report data, usually in PDF or TXT format, cannot be extracted by searching.
@@
The prompt used for LLMs is detailed in Appendix D. We incorporate three specific strategies to alleviate the hallucination issue in LLMs and enhance preprocessing quality: (1) A concise introduction of the text. (2) Mandating the LLM to reference the data source, aiding constraint and verification. (3) Permitting the LLM to indicate ‘not available’ when relevance is uncertain, to avoid fabrication. Appendix E provides a showcase of the text before and after processing.
@@
(b) Data preprocessing. Figure 4 provides word cloud visualizations of constructed text data in the health domain, respectively for extracted facts, extracted predictions, and discarded text. Recall that the target variable here is the influenza patients proportion. Highly relevant words such as "pandemic", "vaccine", and "flu" appear more frequently in the extracted facts; research paper-related words such as "edu", "mdpi", and "university" are more common in the discarded text. Besides, the prediction text primarily contains words describing future, such as "will" and "next". These validate the effectiveness of LLM filtering and disentangling. Furthermore, Table 2 presents a comparison of the token count before and after preprocessing. The substantial decrease validates that LLM summarizing improves usability. Appendix F provides the manual verification results on a subset of the data to further validate the effectiveness of preprocessing using LLMs.
@@ 2.4 Considerations for Fairness and Data Release
To consider fairness, we gather data from both the United States and African regions in the Health domain. As depicted in Figure 2, the numerical data of African region exhibits weaker periodicity. Figure 3 shows that the African region has considerably fewer reports compared to the United States. We urge researchers to consider underrepresented groups when conducting multimodal TSA tasks.
@@ 3 Multimodal Time-Series Forecasting Library: MM-TSFlib
In this section, we aim to illustrate the potential benefits of our Time-MMD for multimodal TSA by focusing on time-series forecasting (TSF), a fundamental TSA task. TSF involves predicting future events or trends based on historical time-series data. While most existing TSF methods primarily depend on numerical series, we aim to extend these unimodal TSF methods to multimodality. To achieve this, we contribute both formulating the multimodal TSF problem as well as introducing MM-TSFlib, the first comprehensive multimodal TSF library.
@@ 3.2 Pioneering Solution for Multimodal TSF
Multimodal Integration Framework. We propose a pioneering multimodal integration framework to extend existing unimodal TSF models to their multimodal versions. As illustrated in Figure 6, our framework features an end-to-end pipeline that integrates open-source language models with various TSF models. Numerical and textual series are independently modeled using unimodal TSF models and LLMs models with projection layers. These outputs are then combined using a learnable linear weighting mechanism to produce the final prediction. To reduce computational costs, we freeze the LLM parameters and train only the additional projection layers. We employ pooling layers to address the inconsistent dimensions of textual variables. This framework features an end-to-end training manner, with marginal trainable parameters overheads.
Multimodal TSF Library. Building upon the multimodal dataset Time-MMD and integration framework, we present the first multimodal TSF library, named MM-TSFlib. MM-TSFlib supports multimodal extensions of over 20 unimodal TSF algorithms through 7 open-source (large) language models, including BERT(11), GPT-2(32) (Small, Medium, Large, Extra-Large), Llama-2-7B(37), and Llama-3-8B333https://llama.meta.com/llama3. We detail the implementations and language models in Appendix G.
@@ 4 Experiments for Multimodal TSF
Based on the constructed MM-TSFlib, we further conduct comprehensive experiments to demonstrate the superiority of multimodal TSF and the high quality of Time-MMD.
@@ 4.1 Experimental Setup
We adhere to the general setups following existing TSF literatures(45; 44; 30). Regarding the horizon window length, we consider a wider range from short to long-term TSF tasks, with four different lengths for each dataset according to frequency. We conduct TSF tasks on all 9 domains of Time-MMD. We employ the widely-adopted mean squared error (MSE) as the evaluation metric.
We comprehensively consider 12 advanced unimodal TSF methods across 4 types including: (1) Transformer-based. We deployed two sets of experiments upon TSF models: Transformer (38), Reformer (20), Informer (55), Autoformer (45), Crossformer (54), Non-stationary Transformer (28),FEDformer (57), iTransformer (26). (2) MLP-based: DLinear (51). (3) Agnostic: FiLM (56). (4) LLM-based: Time-LLM(18). Unless otherwise specified, we use GPT-2-Small as the LLM backbone in MM-TSFlib. More details about setup are provided in Appendix H.
unclear correlation between multimodal TSF performance and LLM natural language capabilities
@@ 4.2 Experimental Results
Our experiments aim to investigate the following five aspects.
@@
Effectiveness of multimodal TSF. Figure 6 shows average MSE results for corresponding unimodal and multimodal versions of each TSF backbone. The multimodal versions consistently outperform corresponding unimodal versions, reducing the mean squared error by over 15% in average and up to 40% in domains with rich textual data. Such significant improvements fully validate the superiority of multimodal TSF and our multimodal framework in Section 3.
@@
Additionally, we observe that different TSF backbones benefit from multimodal extension to varying degrees. For example, the originally inferior Informer exhibits strong multimodal performance, which we attribute to its intrinsic design for modeling long-range dependencies that may benefit more from textual cues. We hope these results inspire more advanced multimodal TSF models.
@@
Quality of Time-MMD dataset. Figure 6 shows that SOTA unimodal TSF models, such as iTransformer and PatchTST, maintain leading unimodal performance, validating the quality of Time-MMD’s numerical data. Moreover, multimodal extension consistently and significantly improves performance by incorporating textual data, confirming the quality of Time-MMD’s textual data
@@
Influence of data domains. Figure 7a shows the relationship between the relevant fact count and the reduced MSE via multimodal extension for each domain. The scatter plot generally illustrates a positive linear correlation, aligning with the innovation of integrating textual information. Besides, domain characteristics also influence multimodal performance, even with a similar fact count. For example, the security domain, focusing on disasters and emergency grants, exhibits higher unpredictability in the future thus benefits less from the historical textual information. This observation highlights the importance of Time-MMD’s coverage of 9 domains.
Influence of the LLM backbone. We varied the LLM backbone in MM-TSFlib and evaluated corresponding multimodal performance on health domain. As shown in Figure 7b, the choice of LLM backbone does not exhibit a significant correlation with multimodal TSF performance. For the GPT2 series, the scaling law is unclear for multimodal TSF, indicating no clear positive correlation between the parameter scale and TSF performance. Across different LLMs, multimodal TSF performance is relatively similar, even between the advanced Llama-3-8B and the earlier BERT. There might be two possible reasons: (1) Our proposed multimodal framework, although effective, still does not fully utilize the power of LLMs, particularly by only fine-tuning through projection layers. (2) Existing LLMs, pre-trained for natural language tasks, may not be directly suitable for multimodal TSF. Either reason suggests that there is significant room for improvement in multimodal TSF.
@@
Influence of the horizon window size. Figure 7c shows the relationship between horizon window size and the average MSE reduction for each domain. Overall, the MSE reduction is stable and promising across different horizon window size, from short term to long term. This demonstrates that the effectiveness of multimodal TSF is robust to different forecast horizon requirements.
5 Broader Impacts

### v4

# Time-MMD: Multi-Domain Multimodal Dataset for Time Series Analysis

@@ Abstract
Time series data are ubiquitous across a wide range of real-world domains. While real-world time series analysis (TSA) requires human experts to integrate numerical series data with multimodal domain-specific knowledge, most existing TSA models rely solely on numerical data, overlooking the significance of information beyond numerical series. This oversight is due to the untapped potential of textual series data and the absence of a comprehensive, high-quality multimodal dataset. To overcome this obstacle, we introduce Time-MMD, the first multi-domain, multimodal time series dataset covering 9 primary data domains. Time-MMD ensures fine-grained modality alignment, eliminates data contamination, and provides high usability. Additionally, we develop MM-TSFlib, the first-cut multimodal time-series forecasting (TSF) library, seamlessly pipelining multimodal TSF evaluations based on Time-MMD for in-depth analyses. Extensive experiments conducted on Time-MMD through MM-TSFlib demonstrate significant performance enhancements by extending unimodal TSF to multimodality, evidenced by over 15% mean squared error reduction in general, and up to 40% in domains with rich textual data. More importantly, our datasets and library revolutionize broader applications, impacts, research topics to advance TSA. The dataset is available at https://github.com/AdityaLab/Time-MMD.
@@ 1 Introduction
Time series (TS) data are ubiquitous across a wide range of domains, including economics, urban computing, and epidemiology (52; 56; 29). Analytical tasks on such datasets hence find broad applications in various real-world scenarios such as energy forecasting, traffic planning, and epidemic policy formulation. Human experts typically complete such Time-Series Analysis (TSA) tasks by integrating multiple modalities of time-series data. For instance, epidemiologists combine numerical data on influenza infections with textual domain knowledge, policies, and reports to predict future epidemiological trends. However, most existing TSA models (59; 42; 78; 44; 81; 32; 80; 48) are unimodal, solely using numerical series.
Recently, with the development of Large Language Models (LLMs), the field of TSA is also undergoing an exciting transformative moment with the integration of natural language (73; 34). Existing LLM-based TSA methods incorporate endogenous text derived from numerical series, such as linguistic descriptions of statistical information, which has demonstrated promising benefits (22; 25; 6; 40). However, the potential of exogenous or auxiliary textual signals—such as information on concurrent events and policies that provide additional context to time series—remains untapped. This observation prompts a crucial question for multimodal TSA: Can multimodal TSA models utilize these exogenous textual signals effectively, thereby enhancing current TSA tasks and enabling new applications?
The primary obstacle in addressing this question lies in the absence of a comprehensive, high-quality multimodal TS dataset, as evidenced by three significant gaps: (1) Narrow data domains. Data characteristics and patterns vary between different domains, such as the periodicity of numerical data and the sparsity of textual data. However, current multimodal TS datasets (15; 70; 11; 49; 6) focus solely on stock prediction tasks in the financial domain, which are unable to represent the diverse data domains. (2) Coarse-grained modality alignment. Existing multimodal TS datasets only ensure that the text and numerical data come from the same domain, such as general stock news and the prices of one specific stock. Clearly, an abundance of irrelevant text diminishes the effectiveness of multimodal TSA. (3) Inherent data contamination. Existing multimodal TS datasets overlook two main reasons of data contamination: (1) Textual data often contains predictions. For example, influenza outlook is a regular section in influenza reports. (2) Outdated test set, particularly the textual data, may have been exposed to LLMs, which are pretrained on vast corpuses. For example, the knowledge cutoff for Llama3-70B is December 2023, which is later than the cutoff dates for most existing multimodal TS datasets. These reasons lead to biased evaluations of general or LLM-based TSA models.
@@
Pioneering Multi-Domain Multimodal Time-Series Dataset. We introduce Time-MMD, the first multi-domain multimodal time-series dataset that addresses the aforementioned gaps: (1) encompasses 9 primary data domains. (2) ensures fine-grained modality alignment through meticulously selected data sources and rigorous filtering steps. (3) disentangles facts and predictions from text; ensures all cutoff dates are up to May 2024. To the best of our knowledge, Time-MMD stands as the inaugural high-quality and comprehensive multimodal time-series dataset. We envision Time-MMD offering exciting opportunities to significantly advance time series analysis through multimodal extensions.
Pilot Multimodal Time-Series Forecasting Study. We develop the first-cut multimodal time-series forecasting (TSF) library, MM-TSFlib, piloting multimodal TSA research based on Time-MMD. Our library MM-TSFlib features an end-to-end pipeline with a seamless interface that allows the integration of any open-source language models with arbitrary TSF models, thereby enabling multimodal TSF tasks. MM-TSFlib facilitates easy exploration of Time-MMD and supports future advancements in multimodal TSA.
Extensive Evaluations with Significant Improvement. We conducted over 1,000 experiments of multimodal TSF on Time-MMD using MM-TSFlib. The multimodal versions outperformed corresponding unimodal versions in 95% of cases, reducing the mean squared error by an average of over 20% and up to 40% in some domains with rich textual data. This significant and consistent improvement demonstrates the high quality of Time-MMD, the effectiveness of MM-TSFlib, and the superiority of multimodal extensions for TSF.
@@ 2.1 Numerical Series Data Construction
Data Source Selection. We select data sources that are (1) reliable, containing verified knowledge; (2) actively released, allowing for updates with new data; and (3) multi-domain, covering various TSA patterns. Appendix C provides considerations for domain selection. Based on these principles, we choose nine data sources from different domains, as shown in Appendix D. Most sources are from government agencies, with the lowest update frequency being six months.
@@
| Domain | Target | Dimension | Frequency | Number of Samples | Timespan |
| Agriculture | Retail Broiler Composite | 1 | Monthly | 1983 - Present | |
| Climate | Drought Level | 5 | Monthly | 1983 - Present | |
| Economy | International Trade Balance | 3 | Monthly | 1989 - Present | |
| Energy | Gasoline Prices | 9 | Weekly | 1996 - Present | |
| Environment | Air Quality Index | 4 | Daily | 1982 - 2023 | |
| Health | Influenza Patients Proportion | 11 | Weekly | 1997 - Present | |
| Security | Disaster and Emergency Grants | 1 | Monthly | 1999 - Present | |
| Social Good | Unemployment Rate | 1 | Monthly | 1950 - Present | |
| Traffic | Travel Volume | 1 | Monthly | 1980 - Present |
@@ 2.2 Textual Series Data Construction
These two data sources complement each other: report data ensures higher relevance but cannot guarantee all-time coverage, while search results cover all times but are highly redundant; search results aggregate multiple data sources, while report data, usually in PDF or TXT format, cannot be extracted by searching. For frequency alignment, the textual data covers multiple frequencies, with search texts enabling daily precision and reports ranging from weekly to monthly and yearly intervals. Appendix D provides a complete list of keywords and reports source.
@@
The prompt used for LLMs is detailed in Appendix F. We incorporate three specific strategies to alleviate the hallucination issue in LLMs and enhance preprocessing quality: (1) A concise introduction of the text. (2) Mandating the LLM to reference the data source, aiding constraint and verification. (3) Permitting the LLM to indicate ‘not available’ when relevance is uncertain, to avoid fabrication. Appendix G provides a showcase of the text before and after processing.
@@
(b) Data preprocessing. Figure 4 provides word cloud visualizations of constructed text data in the health domain, respectively for extracted facts, extracted predictions, and discarded text. Recall that the target variable here is the influenza patients proportion. Highly relevant words such as "pandemic", "vaccine", and "flu" appear more frequently in the extracted facts; research paper-related words such as "edu", "mdpi", and "university" are more common in the discarded text. Besides, the prediction text primarily contains words describing future, such as "will" and "next". These validate the effectiveness of LLM filtering and disentangling. Furthermore, Table 2 presents a comparison of the token count before and after preprocessing. The substantial decrease validates that LLM summarizing improves usability. Appendix H provides the manual verification results on a subset of the data to further validate the effectiveness of preprocessing using LLMs.
@@ 2.4 Considerations for Fairness and Data Release
To consider fairness, we gather data from both the United States and African regions in the Health domain. As depicted in Figure 2, the numerical data of African region exhibits weaker periodicity. Figure 3 shows that the African region has considerably fewer reports compared to the United States. We encourage researchers to consider underrepresented groups when conducting multimodal TSA tasks.
@@ 3 Multimodal Time-Series Forecasting Library: MM-TSFlib
In this section, we aim to illustrate the potential benefits of our Time-MMD for multimodal TSA by focusing on time-series forecasting (TSF), a fundamental TSA task. TSF involves predicting future events or trends based on historical time-series data. While most existing TSF methods primarily depend on numerical series, we aim to extend these unimodal TSF methods to multimodality. To achieve this, we contribute both formulating the multimodal TSF problem as well as introducing MM-TSFlib, the first-cut reasonable approach for multimodal TSF.
@@ 3.2 Pioneering Solution for Multimodal TSF
Multimodal Integration Framework. We propose a pioneering multimodal integration framework to extend existing unimodal TSF models to their multimodal versions. As illustrated in Figure 6, our framework features an end-to-end pipeline that integrates open-source language models with various TSF models. Numerical and textual series are independently modeled using unimodal TSF models and LLMs with projection layers. These outputs are then combined using a learnable linear weighting mechanism to produce the final prediction. To reduce computational costs, we freeze the LLM parameters and train only the additional projection layers. We employ pooling layers to address the inconsistent dimensions of textual variables. This framework features an end-to-end training manner, with minimal overhead in trainable parameters.
Multimodal TSF Library. Building upon the multimodal dataset Time-MMD and integration framework, we present the first multimodal TSF library, named MM-TSFlib. MM-TSFlib supports multimodal extensions of over 20 unimodal TSF algorithms through 7 open-source (large) language models, including BERT (14), GPT-2 (50) (Small, Medium, Large, Extra-Large), Llama-2-7B (58), and Llama-3-8B333https://llama.meta.com/llama3. We detail the implementations and language models in Appendix J.
@@ 4 Experiments for Multimodal TSF
Based on the constructed MM-TSFlib, we further conduct comprehensive experiments to demonstrate the superiority of multimodal TSF and the high quality of Time-MMD."We further investigate the impact of data domains, horizon window size and the text modeling method.
@@ 4.1 Experimental Setup
We adhere to the general setups following existing TSF literatures (67; 66; 48). Regarding the horizon window length, we consider a wider range from short- to long-term TSF tasks, with four different lengths for each dataset according to frequency. We conduct TSF tasks on all 9 domains of Time-MMD. We employ the widely-adopted mean squared error (MSE) as the evaluation metric. A higher MSE indicates a better performance.
We comprehensively consider 12 advanced unimodal TSF methods across 4 types including: (1) Transformer-based: Transformer (59), Reformer (32), Informer (80), Autoformer (67), Crossformer (78), Non-stationary Transformer (44),FEDformer (82), iTransformer (42). (2) MLP-based: DLinear (74). (3) Agnostic: FiLM (81). (4) LLM-based: Time-LLM (25). Unless otherwise specified, we use GPT-2-Small as the LLM backbone in MM-TSFlib. We deployed two sets of experiments upon each TSF model, i.e., both unimodal and multimodal versions. More details about experimental setup are provided in Appendix K.
@@ 4.2 Experimental Results
Our experiments aim to investigate the following five aspects. Appendix O provides detailed results.
@@
Effectiveness of multimodal TSF. Figure 6 shows average MSE results for corresponding unimodal and multimodal versions of each TSF backbone. The multimodal versions consistently outperform corresponding unimodal versions. As detailed in Appendix O, the multimodal versions outperformed their unimodal versions in 95% of over 1,000 experiments, reducing the mean squared error by over 15% in average and up to 40% in domains with rich textual data. Such consistent improvements fully validate the superiority of multimodal TSF and the effectiveness of our proposed multimodal framework in Section 3.
@@
Furthermore, we observe that different TSF backbones benefit from multimodal extension to varying degrees. For example, the originally inferior Informer exhibits strong multimodal performance, which we attribute to its intrinsic design for modeling long-range dependencies that may benefit more from textual cues.
Additional experiments on text integration approach approaches are provided in Appendix L. We hope these results inspire more advanced multimodal TSF solutions.
@@
Quality of Time-MMD dataset. Figure 6 shows that SOTA unimodal TSF models, such as iTransformer and PatchTST, maintain leading unimodal performance, validating the quality of Time-MMD’s numerical data. Moreover, multimodal extension consistently and significantly improves performance by incorporating textual data, confirming the quality of Time-MMD’s textual data.
@@
Influence of data domains. Figure 7a shows the relationship between the relevant fact count and the reduced MSE via multimodal extension for each domain. The scatter plot generally illustrates a positive linear correlation, aligning with the innovation of integrating textual information. Besides, domain characteristics also influence multimodal performance, even with a similar fact count. For example, the security domain, focusing on disasters and emergency grants, exhibits higher unpredictability in the future thus benefits less from the historical textual information. This observation highlights the importance of Time-MMD’s coverage of 9 diverse domains.
@@
Influence of the horizon window size. Figure 7c shows the relationship between horizon window size and the average MSE reduction for each domain. Overall, the MSE reduction is stable and promising across different horizon window size, from short term to long term. These results demonstrate that the effectiveness of multimodal TSF is robust to different forecast horizon requirements.
Influence of the text modeling method. Firstly, we varied the LLM backbone in MM-TSFlib and evaluated corresponding multimodal performance on health domain. As shown in Figure 7b, the choice of LLM backbone does not exhibit a significant correlation with multimodal TSF performance. For the GPT2 series, the scaling law is unclear for multimodal TSF, indicating no clear positive correlation between the parameter scale and TSF performance. Across different LLMs, multimodal TSF performance is relatively similar, even between the advanced Llama-3-8B and the earlier BERT. There might be three possible reasons: (1) Our proposed multimodal framework, although effective, still does not fully utilize the power of LLMs, particularly by only fine-tuning through projection layers. (2) Existing LLMs, pre-trained for natural language tasks, may not be directly suitable for multimodal TSF. (3) The embedding dimension of BERT is 768, much lower than the 4096 dimension size of Llama-3-8B, thereby makes it easier to fine-tune an effective projection layer with limited training data.
Although LLMs are currently the mainstream approach for text modeling, they may not be suitable for certain scenarios, such as unfamiliar domains or unsupported languages. Therefore, we introduce Doc2Vec (33) a text embedding model trained from scratch, as an alternative for text modeling. The experimental results on three datasets across four prediction horizons are provided in Table 6. The results show that Doc2Vec is also effective but generally performs worse than BERT. Doc2Vec is further included in our MM-TSFlib, which will greatly enhance applicability to underrepresented languages and domains.
We provide additional experiments and discussions on multimodal modeling approaches in Appendix N, including the introduction of attention mechanisms (59) and the use of closed-source LLMs, such as GPT-3.5444https://platform.openai.com/docs/models/gpt-3-5-turbo.
In summary, all these observations suggest that there is significant room for improvement in the methodological research for multimodal TSF.
5 Potential Future Works